OpenStudy (anonymous):

Solve: 6x − 10 + 14 = 22

OpenStudy (johnweldon1993):

You want to isolate 'x'....so first....add 10 to both sides... 6x + 14 = 32 now subtract 14 from both sides... 6x = 18 divide both sides by 6 to get x = ...?

OpenStudy (anonymous):

3?

OpenStudy (johnweldon1993):

Correct.....now do you understand the steps I wrote above?

OpenStudy (anonymous):

yes yes i think

OpenStudy (jhannybean):

Woohoo!

OpenStudy (johnweldon1993):

Well I'll go over it 1 more time just for good measure 6x - 10 + 14 = 22 You eventually want to solve for 'x'.....you do this by having x = ...something right? well notice...'x' is all alone on that left side of the equation of what you want...so you want to do this to YOUR equation 6x - 10 + 14 = 22 you can get rid of the -10 on the left side...by adding 10 to it right? But remember what you do to 1 side...you do to the other...so you add 10 to both sides 6x - 10 + 10 + 14 = 22 + 10 6x + 14 = 32 now you do the same thing...but to cancel out +14...you subtract 14...and again..from both sides 6x + 14 - 14 = 32 - 14 6x = 18 and now you want to solve for 'x' not 6x....so you see that x is being multiplied by 6...so to cancel this out...you DIVIDE by 6... because 6x / 6 = 1x...or just x right? so 6x = 18 -------- 6 6 x = 3 Better? :)

OpenStudy (anonymous):

okay yas yas thank you i can answer the rest of the answers

OpenStudy (johnweldon1993):

No problem!